# Entanglement transfer during quantum frequency conversion in gas-filled hollow-core fibers

### v1

@@ Abstract
Quantum transduction is essential for future hybrid quantum networks, connecting devices across different spectral ranges. In this regard, molecular modulation in hollow-core fibers has proven to be exceptional for efficient frequency conversion. However, insights on this conversion method for quantum light have remained elusive beyond standard semiclassical models. This Letter introduces a framework to describe the quantum dynamics of both molecules and photons in agreement with recent experiments and capable of unveiling the ability of molecular modulation to preserve entanglement.
@@
Understanding light-matter interactions at the quantum level lies at the core of the recent developments in quantum technologies Tavis1968 ; Aspelmeyer2014 ; Schmidt2016 ; Blais2021 that are behind sophisticated systems such as the future hybrid quantum networks Kimble2008 . These systems comprise multiple devices such as quantum light sources and memories, fiber transmission lines, etc., which operate across different spectral regions of the optical domain, in sharp contrast to e.g. the microwave superconducting circuits employed in state-of-art quantum computers Girvin2014 ; Gu2017 . Thus, efficient frequency transduction of quantum light states between disparate domains  Li2004 ; Lauk2020 is essential to bridge the operational gaps between nodes Awschalom2021 . While several different approaches to tackle this challenge have been proposed and demonstrated Fan2016 ; Bonsma2022 , molecular modulation in hollow-core anti-resonant fibers (ARFs) filled with gas Russell2014 ; Numkam2023 has recently stood out owing to its near-unity efficiency and exquisite preservation of non-classical correlations Tyumenev2022 . This is facilitated by the tight light-matter confinement in the core Russell2014 , ultralow attenuation over a broad bandwidth Numkam2023 and pressure-adjustable optical properties Bauerschmidt2015 , which make ARFs excellent vehicles for light-based quantum applications Finger2015 ; Antesberger2024 .
@@
On the other hand, in molecular modulation Liang2000 ; Weber2012 at the single-photon level, a quantum light state scatters off the molecular coherence waves pre-excited via stimulated Raman scattering (SRS) in the ARF core, changing its frequency by the appropriate Raman shift without threshold. The corresponding state can be controllably up- or down-converted provided specific phase-matching conditions are fulfilled, which in the case of gas-filled ARFs is achieved by leveraging the fiber dispersion Mridha2019 .
@@
Despite the great potential of ARF-based molecular modulation for quantum transduction Tyumenev2022 , it still remains unclear whether intrinsic quantum properties such as entanglement can be transferred with high fidelity from the original to the target states during the conversion process, a question that cannot be answered using the widely-employed Maxwell-Bloch formalism Raymer1985 ; Raymer1990 . The main reason is its classical treatment of the light fields, although it has been applied to the modelling of certain quantum optical phenomena like photon absorption and emission in weakly-excited atomic clouds Svidzinsky2015 . Therefore, apart from some recent efforts in this direction Wang2023 , a rigorous and accurate description of quantum light-quantum matter interactions in ARF-based frequency conversion down to single-photon limit has, to our best knowledge, so far remained elusive.
@@
In this Letter, we present a quantum Hamiltonian able to describe both the coherence buildup in a molecular gas through SRS, as well as the subsequent thresholdless frequency-conversion process at the single-photon level. In particular, considering the experimental scenario reported in Ref. Tyumenev2022 , we are able to characterize the state of the molecules and predict a complete transfer of entanglement between one of the modes of a Bell state and its corresponding frequency-converted counterpart. Our model can be applied to both confined and free-space geometries and reduces to the semiclassical Maxwell-Bloch formalism Raymer1990 for strong light fields. We expect this general theoretical framework to aid the design, optimization and interpretation of future experiments in light-based quantum technologies using ARFs and their subsequent applications.
@@
The process we want to describe is molecular modulation assisted by SRS Harris1998 ; Nazarkin1999 . We consider two-level molecules in general and, inspired by the experimental results, we focus on the Q$(1)$ vibrational transition of hydrogen as a good two-level approximation. This transition is dipole-forbidden, and therefore needs a two-photon process such as Raman scattering to occur (depicted in Fig. 1).
@@
In this process, the pump photons launched in the fundamental core mode of the H2-filled ARF (illustrated in Fig. 2) are scattered into the Stokes or anti-Stokes frequencies depending on whether they excite or de-excite the molecules, respectively. These transitions are illustrated in Fig. 1, where $\omega_{P}$, $\omega_{S}$, and $\omega_{A}$ are the pump, Stokes, and anti-Stokes angular frequencies, and $\Omega$ represents the Raman shift, such that $\omega_{S}=\omega_{P}-\Omega$ and $\omega_{A}=\omega_{P}+\Omega$. For the H2 gas case, $\Omega/2\pi\approx 125\text{ THz}$ Veirs1987 , the largest molecular shift in nature. Without loss of generality, the light is linearly polarized in our analysis, and therefore rotational states are highly disfavored. Furthermore, hereafter we will consider all the optical frequencies involved in the dynamics contained in the fundamental transmission band of the ARF, i.e. spectrally away from loss-inducing resonances with modes localized in the cladding elements Numkam2023 .
The quantum Hamiltonian describing the pump, Stokes, and anti-Stokes modes of the electric field interacting with a single molecule is expressed as $H=H_{0}+V$ Scully1997 ; Begzjav2021 . On the one hand, we have the unperturbed part of the Hamiltonian $H_{0}$, defined as
@@
where $g_{i,j}^{l}$ is the coupling strength between levels $|i\rangle$ and $|j\rangle$ via bosonic mode $l$, and $\beta_{l}$ is the propagation constant for mode $l$. Let us now eliminate the higher energy levels, i.e. levels with $i>1$, to obtain an effective Hamiltonian describing the interaction of the pump, Stokes, anti-Stokes modes, with the molecule. In order to do this, we go to an interaction picture with respect to $H_{0}$ and perform a rotating-wave approximation, keeping only the static terms up to second order in the coupling strength. That is, by assuming that $g_{i,j}^{l}\sqrt{n_{l}}\ll\hbar|\omega_{i}-\omega_{j}\pm\omega_{l}|$, with $n_{l}$ the number of photons in the mode $l$, we keep only the resonant terms. The only resonances we can identify in this system are $\omega_{P}-\omega_{S}=\Omega=\omega_{A}-\omega_{P}$, with $\Omega\equiv\omega_{1}-\omega_{0}$. Extending this approach to a system with $N$ molecules, we find the effective Hamiltonian SM
@@
where $G_{S(A)}$ is the interaction strength between pump, (anti-)Stokes, and the molecules, $\Delta_{l}^{\pm}$ represent the Stark shifts, and $\Delta\beta\equiv\beta_{P}-\beta_{S}$ and $\Delta\beta^{\prime}\equiv\beta_{A}-\beta_{P}$. The global spin operators are defined through the $1/2$-spin operators as $J_{z}=\oplus_{l=1}^{N}\sigma^{z}_{l}/2$ and $J^{\pm}=\oplus_{l=1}^{N}\sigma^{\pm}_{l}$, with $\sigma^{z}=|1\rangle\langle 1|-|0\rangle\langle 0|$, $\sigma^{+}=|1\rangle\langle 0|$, and $\sigma^{-}=|0\rangle\langle 1|$. They satisfy $[J_{z},J^{\pm}]=\pm J^{\pm}$ and $[J^{+},J^{-}]=2J_{z}$ and, as operators, they act on global spin states of the form $\left|{\bf N/2},m_{z}\right\rangle$, with $m_{z}\in\{-N/2,\ldots,N/2\}$. Note that these operators treat the molecules as an ensemble of two-level systems and they are not representing actual angular momentum of the molecules or light polarization.
In Ref. Tyumenev2022 , the excitation of molecular coherence, that manifests itself as a synchronous oscillation of the gaseous core (see Fig. 2), was achieved using nanosecond-long pump pulses with 115 µJ energy. This means that the initial state of the pump can be considered as a coherent state with $\alpha_{P}\approx 2.48\times 10^{7}$, enabling a few approximations that simplify Eq. (3).
@@
Figure 2: Schematic representation of the experimental layout considered. A pump beam generates Raman vibrational excitations in the gas molecules, preparing them in a coherent and synchronized vibrational motion. During this process, the pump is depleted into the Stokes frequency, as depicted in the figure. The mixing signal simultaneously propagating with the pump perceives the molecular coherence wave and it is scattered to a higher frequency.
@@
Firstly, the annihilation of pump photons leads to the generation of intense laser radiation in the Stokes mode along the fiber length Benabid2002 . Therefore, we may consider a semiclassical approximation, replacing the operators in both the pump and the Stokes modes by classical variables $a_{P}\rightarrow\alpha_{P}$ and $a_{S}\rightarrow\alpha_{S}$ in Eq. (3). Additionally, since the majority of molecules remain in their ground state, the population of the anti-Stokes mode is usually negligible in this process, so we will discard it in our treatment. Finally, we consider that the Stark shifts are also negligible, $\Delta_{P}^{\pm}=\Delta_{S}^{\pm}\approx 0$. Before performing the semiclassical approximation, in order to avoid oscillations with $\Omega$ in the expectation values, we transform the Hamiltonian into an interaction picture with respect to $\Omega J_{z}+\omega_{P}a_{P}^{\dagger}a_{P}+\omega_{S}a_{S}^{\dagger}a_{S}$, obtaining
@@
By evolving the initial state of the molecules $\left|{\bf\frac{N}{2}},-\frac{N}{2}+n\right\rangle$, which corresponds to all molecules in the ground state, under this Hamiltonian for a time $t$, we find the state in the interaction picture SM
@@
Here, we have $s=e^{i\left(\Delta\beta z-\frac{\pi}{2}\right)}\tan\left(G_{S}\alpha_{P}\alpha_{S}t\right)$, where we have assumed that $G_{S}$, $\alpha_{P}$, and $\alpha_{S}$ are real. Interestingly, this state has the functional form of a spin coherent or Bloch state Arecchi1972 .
@@
The emergence of vibrational coherence in the molecular gas, highlighted in green inside the fiber at Fig. 2, originates from the beating between the pump and Stokes fields Mridha2019 ; Raymer1990 . As the amplitude of the coherence wave rises, the pump starts to suffer depletion and the Stokes starts to be amplified. As the depletion continues, the beating between the fields becomes weaker, preventing the generation of new coherence. Meanwhile, the existing coherence wave fades away due to collisional dephasing on a time scale $T_{2}$. Hence, the excited molecular coherence is harvested, within its lifetime, for frequency conversion of an arbitrary mixing signal. Unlike SRS, this frequency conversion process is thresholdless and hence, it can be applied to a single photon. Additionally, the phase-matching conditions governing the feasibility of frequency conversion of the mixing signal are highly influenced by the dispersion contributions from both the gas and the geometry of the waveguide Mridha2019 . In a nutshell, if the difference in the propagation constants of the ARF modes at the original and converted photon frequencies matches the propagation constant of the molecular coherence wave (given by the difference between those of the pump and Stokes fields), energy will be efficiently exchanged during the scattering event, resulting in a modification of the photon frequency according to the molecular Raman shift.
@@
In the following, let us use the developed model to analyze the process of frequency conversion in one of the frequency modes of a maximally-entangled Bell state representing the mixing signal. Indeed, we will convert the mixing mode by launching it simultaneously with the pump beam, while the idler mode remains unperturbed outside of the fiber, and observe the entanglement dynamics between the idler and the mixing and up-converted modes. Considering the experimental conditions of Ref. Tyumenev2022 , the system undergoes a phase-matched transition between a mixing frequency of $\approx 210$ THz (1425 nm in wavelength) and an up-converted frequency of $\approx 335$ THz (895 nm). Even though this is not strictly a Raman transition, it can be described using a Hamiltonian with the same structure as in Eq. (3). This time, the terms describing the mixing to up-converted interaction mimic those of the pump to anti-Stokes, but with the appropriate parameters.
@@
While the idler and the mixing frequency modes are initially prepared in a Bell state, the up-converted mode is not populated; therefore, the initial state is simply $(|0,0,0\rangle+|1,1,0\rangle)/\sqrt{2}$. This notation represents, from left to right and separated by commas, the number of photons in the idler, mixing, and up-converted modes, respectively. Since we typically have a large number of molecules ($\sim 10^{18}$), the state of the molecular ensemble will not change significantly due to the introduction of a single excitation into the system. Thus, in the Hamiltonian characterizing the frequency-conversion process, we replace the global spin operators by their expectation values over the spin coherent state in Eq. (5). That is, we replace $J^{+}\rightarrow\xi^{*}$ and $J^{-}\rightarrow\xi$, with $\xi=\langle s|J^{-}|s\rangle=(N/2)\exp[i\left(\Delta\beta z-\pi/2\right)]\sin\left(2G_{S}\alpha_{P}\alpha_{S}t\right)$. Before, let us transform the Hamiltonian into an interaction picture again with respect to $\Omega J_{z}+\omega_{M}a_{M}^{\dagger}a_{M}+\omega_{U}a_{U}^{\dagger}a_{U}$. Then, the resulting Hamiltonian is
@@
where $G_{U}$ represents the interaction strength between the mixing, the up-converted, and the molecules, and is supposed to be real. Assuming resonance and phase-matching conditions, that are experimentally achievable, we can write $\omega_{U}-\omega_{M}=\Omega$ and $\beta_{U}-\beta_{M}=\Delta\beta$, respectively. The Hamiltonian in Eq. (Entanglement transfer during quantum frequency conversion in gas-filled hollow-core fibers) describes beamsplitter-like dynamics with a time-dependent reflectivity $\sin^{2}G_{U}|\xi|t$, and thus the evolution of a single-photon frequency state can easily be computed. In this framework, the resulting equations for the evolution of the mixing, $N_{M}$, and the up-converted, $N_{U}$, photon numbers are as follows:
$N_{M}$ $=$ $\frac{1}{2}\cos^{2}G_{U}|\xi|t,$ (7)
$N_{U}$ $=$ $\frac{1}{2}\sin^{2}G_{U}|\xi|t.$ (8)
Meanwhile, we also study the dynamics of entanglement between the mixing and the up-converted frequency modes through the concurrence Wootters1998 , an entanglement monotone used for bipartite mixed states. This is an appropriate choice in this case since we have entanglement between the idler and the mixing modes, but also between the idler and up-converted modes. Furthermore, the concurrence completely characterizes the entanglement of formation Hill1997 for a pair of two-level systems. In our case, we find that the idler-mixing and the idler-up-converted concurrences are SM
$\mathcal{C}_{\text{I-M}}$ $=$ $\Big{|}\cos G_{U}|\xi|t\Big{|},$ (9)
$\mathcal{C}_{\text{I-U}}$ $=$ $\Big{|}\sin G_{U}|\xi|t\Big{|}.$ (10)
Notice how entanglement transfer between mixing and up-converted modes is closely related to the evolution of the number of photons in each mode, such that frequency conversion leads to entanglement transfer. The time parameter $t$ here is considered to be related to the propagation distance $z$ inside the fiber used in Ref. Tyumenev2022 , $z=ct$. This is used to represent the evolution of the photon numbers and the concurrences in Fig. 3, where the explicit time dependence of $\alpha_{P}$ and $\alpha_{S}$ has been considered SM . These coefficients are obtained by numerically solving the semiclassical Maxwell-Bloch equations of motion Raymer1990 for the pump and Stokes electric fields, and include a phenomenological factor accounting for the temporal dephasing of the molecular coherence. In addition, in order to obtain the coupling strength, we have estimated the quantization volume based on the geometric properties of the waveguide and the temporal length of the interaction.

### v2

@@ Abstract
Quantum transduction is essential for the future hybrid quantum networks, connecting devices across different spectral ranges. In this regard, molecular modulation in hollow-core fibers has proven to be exceptional for efficient and tunable frequency conversion of arbitrary light fields down to the single-photon limit. However, insights on this conversion method for quantum light have remained elusive beyond standard semiclassical models. In this Letter, we employ a quantum Hamiltonian framework to characterize the behavior of entanglement during molecular modulation, while describing the quantum dynamics of both molecules and photons in agreement with recent experiments. In particular, apart from obtaining analytical expressions for the final opto-molecular states, our model predicts a close correlation between the evolution of the average photon numbers and the transfer of entanglement between the interacting parties. These results will contribute to the development of new fiber-based strategies to tackle the challenges associated with the upcoming generation of lightwave quantum technologies.
pacs:
42.50.-p, 42.65.Ky, 03.65.Ud, 33.80.-b
I Introduction
@@
Understanding light-matter interactions at the quantum level lies at the core of the recent developments in quantum technologies Tavis1968 ; Aspelmeyer2014 ; Schmidt2016 ; Blais2021 that are behind sophisticated systems such as the future hybrid quantum networks Kimble2008 . These systems comprise multiple devices such as quantum light sourcesArakawa2020 ; Zhang2024 , repeatersMunro2015 ; Azuma2023 , memories Heshami2016 ; Lei2023 , fiber transmission lines Wang2022 , etc., which operate across different spectral regions of the optical domain, in sharp contrast to e.g. the microwave superconducting circuits employed in state-of-art quantum computers Girvin2014 ; Gu2017 . Thus, efficient frequency transduction of quantum light states between disparate domains  Li2004 ; Radnaev2010 ; Lauk2020 is essential to bridge the operational gaps between nodes Awschalom2021 . This has encouraged the proposal and demonstration of many different strategies to tackle this challenge in different platforms Gnauck2006 ; Gogyan2008 ; McGuinness2011 ; Ramelow2012 ; Clark2013 ; Williamson2014 ; Donvalkar2014 ; Lefrancois2015 ; Bell2016 ; Fan2016 ; Clemmen2018 ; Bonsma2022 . Recently, molecular modulation in hollow-core anti-resonant fibers (ARFs) filled with gas Russell2014 ; Numkam2023 has stood out owing to its tunability, large frequency shifts, near-unity efficiency and exquisite preservation of non-classical correlations Sokolov2022 ; Tyumenev2022 ; Hamer2024fiber . This is facilitated by the tight light-matter confinement in the core Russell2014 , ultralow attenuation over a broad bandwidth Numkam2023 and pressure-adjustable optical properties Bauerschmidt2015 , which make ARFs excellent vehicles for light-based quantum applications Finger2015 ; Antesberger2024 .
@@
On the other hand, in molecular modulation Liang2000 ; Weber2012 at the single-photon level, a quantum light state scatters off the molecular coherence waves pre-excited via stimulated Raman scattering (SRS) in the ARF core, changing its frequency by the appropriate Raman shift without threshold. The corresponding state can be controllably up- or down-converted provided specific phase-matching conditions are fulfilled, which in the case of gas-filled ARFs is achieved by leveraging the fiber dispersion Hosseini2017 ; Mridha2019 .
@@
Despite the great potential of ARF-based molecular modulation for quantum transduction Tyumenev2022 , it still remains unclear whether intrinsic quantum properties such as entanglement can be transferred with high fidelity from the original to the target states during the conversion process, a question that cannot be answered using the widely-employed Maxwell-Bloch formalism Raymer1985 ; Raymer1990 . The main reason is its classical treatment of the light fields, although it has been applied to the modelling of certain quantum optical phenomena like photon absorption and emission in weakly-excited atomic clouds Svidzinsky2015 . Recent efforts in this direction have provided further insights on the changes in photonic correlations after frequency conversion Wang2023 . Nevertheless, a more detailed description of the internal quantum light-quantum matter interactions down to single-photon limit has, to our best knowledge, so far not been adapted to molecular modulation-based frequency conversion in ARFs.
@@
In this Letter, we describe both the preparation of the molecular state through SRS, as well as the subsequent thresholdless frequency-conversion process at the single-photon level by employing a quantum Hamiltonian model that allows us to study the behavior of entanglement during molecular modulation. In particular, considering the experimental scenario reported in Ref. Tyumenev2022, , we are able to characterize the state of the molecules and predict a complete transfer of entanglement between one of the frequencies of a Bell state and its corresponding frequency-converted counterpart. Our results may aid the design, optimization and interpretation of future experiments in light-based quantum technologies using ARFs and their subsequent applications.
II Quantum description of molecular modulation
II.1 Preparation of the molecular state
@@
The process we want to describe is molecular modulation assisted by SRS Harris1998 ; Nazarkin1999 . We consider two-level molecules in general and, inspired by the recent experiments Tyumenev2022 ; Aghababaei2023 ; Hamer2024 ; Hamer2024fiber , we focus on the $Q(1)$ vibrational transition of hydrogen as a good two-level approximation. This transition is dipole-forbidden, and therefore needs a two-photon process such as Raman scattering to occur (depicted in Fig. 1).
@@
In this process, the pump photons launched in the fundamental core mode of the H2-filled ARF (illustrated in Fig. 2) are scattered into the Stokes or anti-Stokes frequencies depending on whether they excite or de-excite the molecules, respectively. These transitions are illustrated in Fig. 1, where $\omega_{P}$, $\omega_{S}$, and $\omega_{A}$ are the central angular frequencies of the narrowband pump, Stokes, and anti-Stokes signals, and $\Omega$ represents the Raman shift, such that $\omega_{S}=\omega_{P}-\Omega$ and $\omega_{A}=\omega_{P}+\Omega$. For the H2 gas case, $\Omega/2\pi\approx 125\text{ THz}$ Veirs1987 , the largest molecular shift in nature. Without loss of generality, the light is linearly polarized in our analysis, and therefore rotational states are highly disfavored. Furthermore, hereafter we will consider all the optical frequencies involved in the dynamics contained in the fundamental transmission band of the ARF, i.e. spectrally away from loss-inducing resonances with spatial modes localized in the cladding elements Numkam2023 .
The quantum Hamiltonian describing the pump, Stokes, and anti-Stokes frequencies of the electric field interacting with a single molecule is expressed as $H=H_{0}+V$ Scully1997 ; Begzjav2022 . On the one hand, we have the unperturbed part of the Hamiltonian $H_{0}$, defined as
@@
where $g_{i,j}^{l}$ is the coupling strength between levels $|i\rangle$ and $|j\rangle$ via the bosonic mode $l$, and $\beta_{l}$ is the propagation constant for frequency $\omega_{l}$. Let us now eliminate the higher energy levels, i.e. levels with $i>1$, to obtain an effective Hamiltonian describing the interaction of the pump, Stokes, anti-Stokes frequencies, with the molecule. In order to do this, we go to an interaction picture with respect to $H_{0}$ and perform a rotating-wave approximation, keeping only the static terms up to second order in the coupling strength. That is, by assuming that $g_{i,j}^{l}\sqrt{N_{l}}\ll\hbar|\omega_{i}-\omega_{j}\pm\omega_{l}|$, with $N_{l}$ the number of photons with frequency $\omega_{l}$, we keep only the resonant terms. The relevant resonances in this system are $\omega_{P}-\omega_{S}=\Omega=\omega_{A}-\omega_{P}$, with $\Omega\equiv\omega_{1}-\omega_{0}$. Extending this approach to a system with $N$ molecules results in the following effective Hamiltonian Walls1971 :
@@
where $G_{S(A)}$ is the interaction strength between pump, (anti-) Stokes, and the molecules, $\Delta_{l}^{\pm}$ represent the Stark shifts, and $\Delta\beta\equiv\beta_{P}-\beta_{S}$ and $\Delta\beta^{\prime}\equiv\beta_{A}-\beta_{P}$. The global spin operators are defined through the $1/2$-spin operators as $J_{z}=\oplus_{l=1}^{N}\sigma^{z}_{l}/2$ and $J^{\pm}=\oplus_{l=1}^{N}\sigma^{\pm}_{l}$, with $\sigma^{z}=|1\rangle\langle 1|-|0\rangle\langle 0|$, $\sigma^{+}=|1\rangle\langle 0|$, and $\sigma^{-}=|0\rangle\langle 1|$. They satisfy $[J_{z},J^{\pm}]=\pm J^{\pm}$ and $[J^{+},J^{-}]=2J_{z}$ and, as operators, they act on global spin states of the form $\left|{\bf N/2},m_{z}\right\rangle$, with $m_{z}\in\{-N/2,\ldots,N/2\}$. Note that these operators treat the molecular gas as an ensemble of two-level systems and they are not representing actual angular momentum of the molecules or light polarization. More details on the derivation of $H_{\text{eff}}$ can be found in the supplementary material.
The excitation of molecular coherence manifests itself as a synchronous oscillation of the gaseous core. This is shown in Fig. 2, where the molecules are depicted equidistantly along a longitudinal axis with the same orientation for illustrative purposes, whereas in reality the gas fills the whole interior of the fiber and the molecules are randomly oriented. In this regard, the interaction will be more significant with the percentage of molecules aligned with the linearly polarized fields, which is already taken into account in the coupling constants obtained phenomenologically (see supplementary material). In Ref. Tyumenev2022, , the molecular coherence was generated by a nanosecond pump pulse with an energy of 115 µJ. This means that the initial state of the pump can be considered as a coherent state with $\alpha_{P}\approx 2.48\times 10^{7}$, enabling a few approximations that simplify Eq. (3).
@@
Figure 2: Schematic representation of the experimental layout considered. A pump beam generates Raman vibrational excitations in the gas molecules, preparing them in a coherent and synchronized vibrational motion. During this process, the pump is depleted into the Stokes frequency, as represented in the figure. The mixing signal simultaneously propagating with the pump perceives the molecular coherence wave and it is scattered to a higher frequency. Although the gas molecules are depicted equidistantly along a longitudinal axis for illustrative purposes, they actually fill the whole interior of the fiber and are randomly located and oriented.
@@
Firstly, the annihilation of pump photons leads to the generation of intense laser radiation at the down-shifted Stokes frequency along the fiber length Benabid2002 . Therefore, we may consider a semiclassical approximation, replacing the operators in both pump and Stokes frequencies by classical variables $a_{P}\rightarrow\alpha_{P}$ and $a_{S}\rightarrow\alpha_{S}$ in Eq. (3). Additionally, since the majority of molecules remain in their ground state, the photon population of the anti-Stokes frequency is usually negligible in this process, allowing us to discard it in our analysis. Finally, we consider that the Stark shifts are also negligible, $\Delta_{P}^{\pm}=\Delta_{S}^{\pm}\approx 0$. Before performing the semiclassical approximation, in order to avoid oscillations with $\Omega$ in the expectation values, we transform the Hamiltonian into an interaction picture with respect to $\Omega J_{z}+\omega_{P}a_{P}^{\dagger}a_{P}+\omega_{S}a_{S}^{\dagger}a_{S}$, obtaining
@@
By evolving the initial state of the molecules $\left|{\bf\frac{N}{2}},-\frac{N}{2}\right\rangle$, which corresponds to all molecules in the ground state, under this Hamiltonian for a time $t$, we find the state in the interaction picture
@@
Here, we have
$s=-ie^{i\Delta\beta z}\tan\left(G_{S}\alpha_{P}\alpha_{S}t\right),$ (6)
where we have assumed that $G_{S}$, $\alpha_{P}$, and $\alpha_{S}$ are real. The supplementary material and Ref. Arecchi1972, include the mathematical steps to find $|s\rangle$.
@@
The emergence of vibrational coherence in the molecular gas, highlighted in green inside the fiber at Fig. 2, originates from the beating between the pump and Stokes fields Mridha2019 ; Raymer1990 ; Chen2024 . As the amplitude of the coherence wave rises, the pump starts to suffer depletion and the Stokes starts to be amplified. As the depletion continues, the beating between the fields becomes weaker, preventing the generation of new coherence. Meanwhile, the existing coherence wave fades away due to collisional dephasing on a time scale $T_{2}$. However, in a temporal reference frame moving with the co-propagating pump and Stokes pulses at their group velocity, new coherence will be found over a longer time scale as light travels through the fiber, since the amplitude of the coherence wave at a given relative time coordinate is the result of the integrated effects induced by these pulses at previous times Raymer1990 . Hence, the excited molecular coherence is harvested, within its lifetime, for frequency conversion of an arbitrary mixing signal. Unlike usual SRS, in which the power of the scattered pump beam needs to be higher than a minimum threshold, this frequency conversion process is thresholdless and hence, it can be applied to a single photon. Additionally, the phase-matching conditions governing the feasibility of frequency conversion of the mixing signal are highly influenced by the dispersion contributions from both the gas and the geometry of the waveguide Hosseini2017 ; Mridha2019 . In a nutshell, if the difference in the propagation constants of the ARF modes at the original and converted photon frequencies matches the propagation constant of the molecular coherence wave (given by the difference between those of the pump and Stokes fields), energy will be efficiently exchanged during the scattering event, resulting in a modification of the photon frequency according to the molecular Raman shift.
II.2 Mixing frequency conversion
@@
In the following, let us use the developed framework to analyze the frequency conversion process of one of the frequency components of a maximally-entangled Bell state. Indeed, we will convert the mixing frequency by launching it simultaneously with the pump beam, while the idler frequency remains unperturbed outside of the fiber, and observe the entanglement dynamics between the idler and the mixing and up-converted frequencies. Considering the conditions of Ref. Tyumenev2022, , the experimental system undergoes a phase-matched transition between a mixing frequency of $\approx 210$ THz (1425 nm in wavelength) and an up-converted frequency of $\approx 335$ THz (895 nm). Even though this is not strictly a Raman transition, it can be described using a Hamiltonian with the same structure as in Eq. (3). This time, the terms describing the mixing to up-converted interaction mimic those of the pump to anti-Stokes, but with the appropriate parameters.
@@
While the idler and the mixing fields are initially prepared in a Bell state, the up-converted field is in the vacuum state; therefore, the initial state is simply $(|0,0,0\rangle+|1,1,0\rangle)/\sqrt{2}$. This notation represents, from left to right and separated by commas, the number of photons in the idler, mixing, and up-converted frequencies, respectively. Since in the experimentsTyumenev2022 ; Bauerschmidt2015 ; Mridha2019 we typically have a large number of molecules ($\sim 10^{18}$), the energy state of the molecular ensemble will not change significantly due to the introduction of a single excitation into the system. Thus, in the Hamiltonian characterizing the frequency-conversion process, we replace the global spin operators by their expectation values over the spin coherent state in Eq. (5). That is, we replace $J^{+}\rightarrow\xi^{*}$ and $J^{-}\rightarrow\xi$, with
$\xi=\langle s|J^{-}|s\rangle=-ie^{i\Delta\beta z}\frac{N}{2}\sin\left(2G_{S}\alpha_{P}\alpha_{S}t\right).$ (7)
Before continuing, let us transform the Hamiltonian into an interaction picture again with respect to $\Omega J_{z}+\omega_{M}a_{M}^{\dagger}a_{M}+\omega_{U}a_{U}^{\dagger}a_{U}$, where $\omega_{M}$ and $\omega_{U}$ are the mixing and up-converted angular frequencies. Then, the resulting Hamiltonian is
@@
where $G_{U}$ represents the interaction strength between the mixing, the up-converted, and the molecules, and is supposed to be real. If we assume that, as in the experiments, the phase matching and resonance conditions are satisfied, we would have $\beta_{U}-\beta_{M}=\Delta\beta\equiv\beta_{P}-\beta_{S}$ and $\omega_{U}-\omega_{M}=\Omega\equiv\omega_{P}-\omega_{S}$, respectively. Note that we do not necessarily require $\omega_{l}=\omega_{m}$ or $\beta_{l}=\beta_{m}$ with $l\in\{P,S\}$ and $m\in\{M,U\}$ and that, experimentally, these conditions can be adjusted to carry out frequency down-conversion as well, owing to the excellent tunability of gas-filled ARFs Mridha2019 ; Hosseini2017 . Note that this Hamiltonian is similar to others found in the context of quantum frequency conversion Tucker1969 ; Kumar1990 . Using these conditions and $H_{\xi}^{I}$ to evolve the Bell state considered for the idler and the mixing photons we obtain a state that, when transformed back to the Schrödinger picture, becomes
$|\psi\rangle=\frac{1}{\sqrt{2}}\Big{(}|0,0,0\rangle+\cos\left(G_{U}|\xi|t\right)|1,1,0\rangle\\ +e^{-i\Omega t}\sin\left(G_{U}|\xi|t\right)|1,0,1\rangle\Big{)},$ (9)
which is heavily modulated by the state of the gaseous molecular ensemble via $\xi$, and is analogous to the ones found in other quantum frequency conversion schemes Clark2013 .